OpenStudy (anonymous):

A drug trial is testing the effectiveness of two drugs. If 30 patients are given Drug A, 10 patients are given Drug B, and 90 patients are given a placebo, what is the probability that a patient will not be given a placebo?

OpenStudy (anonymous):

4/13 3/13 9/13 1/13

OpenStudy (anonymous):

it was A

OpenStudy (anonymous):

Total amount of people involved in the trial: 30+10+90=130 People not given Placebo 30+10=40 Probability given placebo is 40/130=4/13

OpenStudy (anonymous):

thanks